OpenStudy (anonymous):

Which of the following is the complete factorization of 10x - 3 - 3x2? (3x - 1)(x + 3) -(3x + 1)(x - 3 )*********** -(3x - 1)(x - 3) I think it is the one with ********

OpenStudy (campbell_st):

i think you're incorrect..

Parth (parthkohli):

\[-3x^2+ 10x - 3 = -3x^2 + x + 9x - 3 = -x(3x -1) + 3(3x - 1) \\ = (-x + 3)(3x - 1) = \cdots\]

OpenStudy (campbell_st):

your choice would have + 3 as the constant when you distribute

Parth (parthkohli):

\[=-1 \times (x -3) \times (3x - 1) = \cdots\]